OpenStudy (anonymous):

Which of the following is the best starting point for approximating the square root of to the nearest whole number? √1<√73<100 √64<√73<√81 √72<√73<√74 √70<√73<√80

OpenStudy (anonymous):

i suppose its the second option because the two outside values are 8 and 9 (i think!) 73 is close to 81 than to 64 so square root is 9 to neares whole number

OpenStudy (anonymous):

one useful way of sqaure two numbers ending in 0.5 is for example 8.5 * 8.5 = 8*9 + 0.25 = 72.25